# A mathematical model for wind-generated particle-fluid flow fields with an application to the helicopter cloud problem

### v1

@@ 1 Introduction
A particular example of such wind-driven particle flow fields is the scenario in which a helicopter may descend towards a bed of sand, the helicopter cloud problem (often refered to as the ‘brownout problem’ in the aero-engineering literature). In this case, the local interaction between the down-draft and swirling flow, generated by the helicopter rotor, and the upper surface of the otherwise static sand bed, entrains sand particles from a thin interfacial layer into the fluidized region, which then flow in the form of a high velocity particle laden cloud around the helicopter body, which we refer to as the helicopter cloud, and a consequence of which is generally a significant deterioration in visibility for the helicopter pilot - when this becomes too severe, it is referred to as brownout. This problem has received much attention in the engineering literature, but most work to date has primarily been driven by observation and experimentation. Attempts to model the physics of brownout include those by Wachspress et al. (2008); Phillips & Brown (2009); Phillips et al. (2011); GLG; Porcù et al. (2020); Tan et al. (2021, 2022a, 2022b); Li et al. (2023); Lin et al. (2024). These modelling approaches broadly split into two classes, which we now delineate.
@@
The first approach is purely computational, and treats each particle within the fluid flow as an individual entity, with no effort being made to address the entrainment or detrainment of particles into the particle laden flow via local interaction of the flow with the upper surface of the otherwise static particle bed. Indeed, the particle bed is removed from consideration, and instead the particles are simply placed into the flow in a random spatial distribution at an initial reference time. Subsequently each particle is tracked as it moves within the fluid flow according to its own dynamical equation of motion under the action of the locally induced fluid interaction forces and gravity (the following, amongst many others, are papers adopting this principal strategy as a key modelling component: GLG; Porcù et al. (2020); Tan et al. (2021, 2022a, 2022b); Li et al. (2023); Lin et al. (2024)). Although this approach does give rational information about particulate behaviour once the particles have made their way into the fluidized flow, the lack of consideration of the entrainment and detrainment process at the interface between the fluidized region and the static particle bed is a very severe drawback for this modelling approach - it is this interaction process that is the fundamental and key process in initiating and driving the whole phenomenon of the particle-cloud. In addition, given that particle clouds in the fluidized region are generally concentrated rather than dilute, with a low fluid voidage (high particle concentration), the approach of treating each particle in the fluidized region as an individual entity is exceptionally inefficient in computational time, and a two phase flow approach is much more natural and efficient.
@@
The second modelling approach is to introduce a continuum particle density field in the fluidized region (measuring particle volume per unit spatial volume of the two phase flow), and then to postulate that this field satisfies a suitable advection-diffusion PDE throughout the flow field. This approach does go some way to address the entrainment and detrainment of particles from the static particle bed. Specifically, this is done in a phenomenological way by the introduction of a localised particle mass source term into the advection-diffusion PDE. This source term is localised in space, so that it acts only in a thin neighbourhood of the interface between the fluidized region and the static particle bed, and is designed to represent the localised input/ouput of particles from the static bed into the fluidized region (see, in particular, Phillips et al. (2011)). The nature of this source term is purely phenomenological, and empirically based on the very particular situation under immediate consideration; as such, the ability of this approach to capture, in general, a decent representation of the key rational mechanism of entrainment/detrainment at the interface of the fluidized region and the static bed is very limited, and depends entirely, and critically, upon the specific choice of this term. This serious limitation is well addressed in the paper of Phillips et al. (2011). In their conclusions, they state that further developments in the theory for the brownout problem depend crucially on an understanding “…particularly of how sediment leaves the ground and enters the flow around the helicopter in the first place …”. Moreover, these authors state (Phillips et al., 2011, p.126) that “…the physics of sediment entrainment from the ground …is no doubt the area which could most benefit from further research and, indeed, where a breakthrough in understanding could contribute most to improving our confidence in the ability of predictive methods to capture the detailed mechanisms at the origin of the brownout cloud …”. The key message is that, although this approach does address local particle entrainment and detrainment from the static particle bed, a thorough understanding and implementation of the fundamental interactive mechanics governing these two principal processes and their coupling with the fluidized region, is essential for improving predictability in the problem.
@@
In this paper we address these issues directly and generically. We consider the general situation of an incompressible fluid overlying an otherwise static particle bed, and allow for the general motion of the fluid to interact with the particle bed in a thin transition layer, in which the local force balance between the fluid flow and the incipiently fluidized particles is fully accounted for on the thickness scale of this thin transition region, which rationally accounts for both the entrainment and detrainment of particles from the static bed and into the fully fluidized region and vice versa. The resolution of the dynamics on the scale of this thin transition layer leads to an additional macroscopic boundary condition on local void fraction which must be applied at the interface of the fully fluidized region and the particle bed surface. This closes the macroscopic two phase flow problem in the fully fluidized region, with now the entrainment and detrainment mechanism rationally accounted for and the need for phenomenological parameters removed. This closed model can now be used to formulate the problem in the fluidized region for any particular circumstances under consideration. To illuminate this theoretical model, we apply it in detail to a simple version of the helicopter cloud problem in the latter part of the paper.
An outline of the paper is as follows. In §2 we give a detailed description of the physical model, and in particular our approach to rationally modelling the dynamics of particle entrainment and detrainment in an interfacial layer. This is followed in §3 by the formulation of the physical model as a mathematical model, with a detailed treatment of the local dynamics in the interfacial layer, and the subsequent development of the corresponding macroscopic boundary conditions which close the full two-phase flow problem in the fully fluidized region. This is followed in  §4 by bringing together these components in the formulation of the full closed mathematical problem in the fluidized region. This mathematical problem is made dimensionless in §5, where the formulation is reduced to the determination of the voidage field and a fluid velocity potential. The application to the helicopter cloud problem is developed in §6, where the effect of the helicopter close to the surface is modelled by a suitably chosen fluid dipole and fluid line-vortex. The boundary value problems which are associated with this application are formulated in a form suitable for numerical solution via finite-difference approximation; in §7 we give precise details of our numerical method, and present numerical examples demonstrating how varying certain parameters may change qualitative features of the voidage field.
@@ 2 The physical model
[…]
and we require the interfacial layer thickness to be small compared to the continuum length scale in the fluidized region, so that
$\delta\ll l.$ (3)
@@
The static particle bed region is taken as a particle store, with the packed particles in static equilibrium (without strong cohesion). For the specific problem addressed in §6, namely the helicopter cloud problem, the continuum length scale $l$ will be taken as the geometric dimension of the helicopter blade length. The hydrodynamic effect of the helicopter is modelled as a fluid dipole, aligned with the axis of blade rotation and located at the helicopter blade rotation point, to represent the fluid down-draft, together with a line-vortex through the axis of rotation of the blades, to represent the fluid swirl.
@@ 3.2 The interfacial layer
which is the Magnus force induced on an individual rolling particle grain as it moves along the interfacial layer. We anticipate that in this incipiently fluidized interfacial layer this Magnus force dominates over the Saffman lift force due to the local shear in the fluid flow. The function $\Phi(E_{I})$ is anticipated to be, in general, a monotone increasing in $E_{I}$, with $\Phi(E_{s})<1$ and $\Phi(1)=1$. Without detailed experimental evidence, we adopt the simple linear form approximation
@@
In (33) the local voidage in the interfacial layer $E_{I}$ is related to the local tangential fluid velocity in the interfacial layer, $\mathbf{u}_{I}$, whilst (32) relates the local tangential particle velocity in the interfacial layer $\mathbf{v}_{I}$ to the local tangential fluid velocity in the interfacial layer $\mathbf{u}_{I}$, and the form is similar to a mobile bed-load sediment transport function (see, for example, Bagnold (1941) for the classical discussion of bed-load sediment transport formula in a wind-blown sand environment). We next consider the motion of the interfacial layer.
@@ 3.3 Interfacial layer motion
Figure 3: Element through static bed and interfacial layer.
@@ 5 The dimensionless problem
$\bar{R}_{p}E(\mathbf{u}_{t}+(\mathbf{u}\cdot\nabla)\mathbf{u})=-\nabla p+GE\mathbf{k}-(1-E)\bar{\beta}(E)(\mathbf{u}-\mathbf{v}),$ (53)
$\rho\bar{R}_{p}(1-E)(\mathbf{v}_{t}+(\mathbf{v}\cdot\nabla)\mathbf{v})=\alpha\nabla E-\rho G(1-E)\mathbf{k}+(1-E)\bar{\beta}(E)(\mathbf{u}-\mathbf{v}),$ (54)
@@
on $z=\xi$, with $\mathbf{r}_{h}\in\mathbb{R}^{2}$, $t>0$. Here we have introduced the dimensionless parameters, $\rho=\rho_{s}/\rho_{f}$, giving the density ratio of the particulate material to the fluid material, whilst
$\bar{R}_{p}=\frac{aR_{p}}{l}$
where $R_{p}=\rho_{f}u_{s}a/\mu$ is the single particle Reynold’s number, and $\bar{R}_{p}$ measures the ratio of the inertia terms to the drag terms in the momentum equations (53) and (54). In addition $G=\rho_{f}ga^{2}/\mu u_{s}$ measures the ratio of the gravity terms to the drag terms in the momentum equations (53) and (54), whilst $\alpha=p_{0}a^{2}/l\mu u_{s}$ measures the ratio of the particle-particle pressure term to the drag term in the momentum equation (54). Finally, $\gamma=3u_{s}^{2}\rho_{f}/ag\rho_{s}$ measures the ratio of the lift force to the gravity force acting on the particles in the interfacial layer. For convenience, we have also written
@@
$\rho\gg 1,\quad\bar{R}_{p}\ll\frac{\alpha}{\rho}\ll 1,\quad G\ll\frac{\alpha}{\rho}\ll 1.$
@@
$(2-E_{s}-E)\nabla\phi\cdot\mathbf{k}+\alpha\nabla H(E)\cdot\mathbf{k}=0,$ (75)
$\xi_{t}=\nabla\phi\cdot\mathbf{k},$ (76)
@@
[…]
@@
[…]
@@ 6 The helicopter cloud problem
[…]
Here $(R,\theta,z)$ are the usual cylindrical polar coordinates in relation to the Cartesian coordinates $(x,y,z)$ which were introduced earlier, with unit vectors $\hat{\mathbf{R}}$, $\hat{\mathbf{\theta}}$ and $\mathbf{k}$. In (83), $\phi_{d}$ represents the harmonic half-space dipole, given by
@@
Here $\hat{\mathbf{N}}$ is the unit outward normal on the cylindrical shell. It remains to consider the far field boundary conditions at large distances from the cylindrical shell surface $\partial C$. We require
@@
[…]
@@
$\nabla\bar{\phi}(R,z)=a(R,z)\hat{\mathbf{R}}+\frac{\Omega}{2\pi R}\hat{\mathbf{\theta}}+b(R,z)\mathbf{k},$ (94)
@@
[…]
@@
[…]
as $\delta\beta_{0}\rightarrow 0$ uniformly for $E\in[E_{s},1]$. On adopting (103), the problem [BVP] finally becomes, in terms of the cylindrical polar coordinates $(R,\theta,z)$,
$E=g(R)\mbox{ on }z=0,\quad R\geq 0,$ (105)
$E_{z}=0\mbox{ on }z=z_{d}^{+},z_{d}^{-},\quad R\in(0,\Delta),$ (109)
Here
for $(R,z)\in\omega$, whilst
$z_{d}^{+}=z_{d}+\frac{\Delta}{2},\quad z_{d}^{-}=z_{d}-\frac{\Delta}{2},$
and
[…]
The domain $\omega$ is illustrated in figure 6.
@@
Before proceeding further with [BVP], we consider the detailed form of $g:[0,\infty)\to\mathbb{R}$, as defined in (99). First we examine $|\nabla_{h}\bar{\phi}(X,0)|^{2}$ for $X\in(0,\infty)$, the form of which depends upon the two parameters $\Omega$ and $z_{d}$. We begin by observing that

### v2

@@ 1 Introduction
A particular example of such wind-driven particle flow fields is the scenario in which a helicopter may descend towards a bed of sand, the helicopter cloud problem (often refered to as the ‘brownout problem’ in the aero-engineering literature). In this case, the local interaction between the down-draft and swirling flow, generated by the helicopter rotor, and the upper surface of the otherwise static sand bed, entrains sand particles from a thin interfacial layer into the fluidized region, which then flow in the form of a high velocity particle laden cloud around the helicopter body, which we refer to as the helicopter cloud, and a consequence of which is generally a significant deterioration in visibility for the helicopter pilot - when this becomes too severe, it is referred to as brownout. This problem has received much attention in the engineering literature, but most work to date has primarily been driven by observation and experimentation. Attempts to model the physics of brownout include those by Wachspress et al. (2008); Phillips & Brown (2009); Phillips et al. (2011); Govindarajan et al. (2013); Porcù et al. (2020); Tan et al. (2021, 2022a, 2022b); Li et al. (2023); Lin et al. (2024). In addition a very recent thesis by Rovere (2022) reviews particle tracking simulations together with experiments and high speed photographic evidence related to the helicopter cloud problem, and provides an excellent survey of the engineering literature in this area. These modelling approaches broadly split into two classes, which we now delineate.
@@
The first approach is purely computational, and treats each particle within the fluid flow as an individual entity, with no effort being made to address the entrainment or detrainment of particles into the particle laden flow via local interaction of the flow with the upper surface of the otherwise static particle bed. Indeed, the particle bed is removed from consideration, and instead the particles are simply placed into the flow in a random spatial distribution at an initial reference time. Subsequently each particle is tracked as it moves within the fluid flow according to its own dynamical equation of motion under the action of the locally induced fluid interaction forces and gravity (the following, amongst many others, are papers adopting this principal strategy as a key modelling component: Govindarajan et al. (2013); Porcù et al. (2020); Tan et al. (2021, 2022a, 2022b); Li et al. (2023); Lin et al. (2024)). Although this approach does give rational information about particulate behaviour once the particles have made their way into the fluidized flow, the lack of consideration of the entrainment and detrainment process at the interface between the fluidized region and the static particle bed is a very severe drawback for this modelling approach - it is this interaction process that is the fundamental and key process in initiating and driving the whole phenomenon of the particle cloud. In addition, given that particle clouds in the fluidized region are generally concentrated rather than dilute, with a low fluid voidage (high particle concentration), the approach of treating each particle in the fluidized region as an individual entity is exceptionally inefficient in computational time, and a two-phase flow approach is much more natural and efficient.
@@
The second modelling approach is to introduce a continuum particle density field in the fluidized region (measuring particle volume per unit spatial volume of the two phase flow), and then to postulate that this field satisfies a suitable advection-diffusion PDE throughout the flow field. This approach does go some way to address the entrainment and detrainment of particles from the static particle bed. Specifically, this is done in a phenomenological way by the introduction of a localised particle mass source term into the advection-diffusion PDE. This source term is localised in space, so that it acts only in a thin neighbourhood of the interface between the fluidized region and the static particle bed, and is designed to represent the localised input/ouput of particles from the static bed into the fluidized region (see, in particular, Phillips et al. (2011)). The nature of this source term is purely phenomenological, and empirically based on the very particular situation under immediate consideration and must be recalibrated in every specific example; as such, the ability of this approach to capture, in general, a decent representation of the key rational mechanism of entrainment/detrainment at the interface of the fluidized region and the static bed is very limited, and depends entirely, and critically, upon the specific choice of this term. This serious limitation is well addressed in the paper of Phillips et al. (2011). In their conclusions, they state that further developments in the theory for the brownout problem depend crucially on an understanding “…particularly of how sediment leaves the ground and enters the flow around the helicopter in the first place …”. Moreover, these authors state (Phillips et al., 2011, p.126) that “…the physics of sediment entrainment from the ground …is no doubt the area which could most benefit from further research and, indeed, where a breakthrough in understanding could contribute most to improving our confidence in the ability of predictive methods to capture the detailed mechanisms at the origin of the brownout cloud …”. The key message is that, although this approach does address local particle entrainment and detrainment from the static particle bed, a thorough understanding and implementation of the fundamental interactive mechanics governing these two principal processes and their coupling with the fluidized region, is essential for improving predictability in the problem. It is this serious defect which we address from fundamental first principles in this paper, which leads to a natural macroscopic boundary condition on void fraction to be applied at the interface of the fluidized region and the static bed, which closes the continuum scale problem two phase flow in the fluidized region.
@@
In this paper we address these issues directly and generically. We consider the general situation of an incompressible fluid overlying an otherwise static particle bed, and allow for the general motion of the fluid to interact with the particle bed in a thin transition layer, in which the local force balance between the fluid flow and the incipiently fluidized particles is fully accounted for on the thickness scale of this thin transition region, which rationally accounts for both the entrainment and detrainment of particles from the static bed and into the fully fluidized region and vice versa. The resolution of the dynamics on the scale of this thin transition layer leads to an additional macroscopic boundary condition on local void fraction which must be applied at the interface of the fully fluidized region and the particle bed surface. This closes the macroscopic two phase flow problem in the fully fluidized region, with now the entrainment and detrainment mechanism rationally accounted for and the need for phenomenological parameters removed. This closed model can now be used to formulate the problem in the fluidized region for any particular circumstances under consideration. To illuminate this theoretical model, we apply it in detail to a simple version of the helicopter cloud problem in the latter part of the paper. Although direct experimental comparisons for the results from our model are not possible at present without a specifically designed experimental program, it is worthwhile referring to the high speed photography in Rovere (2022), through the images labelled as Figure 1.1, Figure 1.2, and, in particular Figure 2.6, which gives a good qualitative comparison with our theoretical examples presented in §7 of this paper.
An outline of the paper is as follows. In §2 we give a detailed description of the physical model, and in particular our approach to rationally modelling the dynamics of particle entrainment and detrainment in an interfacial layer. This is followed in §3 by the formulation of the physical model as a mathematical model, with a detailed treatment of the local dynamics in the interfacial layer, and the subsequent development of the corresponding macroscopic boundary conditions which close the full two-phase flow problem in the fully fluidized region. This is followed in  §4 by bringing together these components in the formulation of the full closed mathematical problem in the fluidized region. This mathematical problem is made dimensionless in §5, where the formulation is reduced to the determination of the voidage field and a fluid velocity potential. The application to the helicopter cloud problem is developed in §6, where the effect of the helicopter close to the surface is modelled by a suitably chosen fluid dipole and fluid line-vortex. The boundary value problems which are associated with this application are formulated in a form suitable for numerical solution via finite-difference approximation; in §7 we present numerical examples demonstrating how varying certain parameters may change qualitative features of the voidage field, with precise details of our numerical method and its accuracy given in Appendix B.
@@ 2 The physical model
$\delta\approx 20a,$ (2)
so that, formally, $a\ll\delta$ and the interfacial layer is thin compared to the two-phase flow continuum length scale $l$, but sufficiently wide to contain enough particles for local average particle velocities to be used. Thus we formally require,
$a\ll\delta\ll l.$ (3)
@@
Connection between the interfacial layer and the bulk fluidized region is made by requiring continuity of the normal particle and fluid mass flow fields and the voidage field at a well defined interface. Of course, this interface is a model device to replace the rapid transition of normal length scale from the interfacial layer into the bulk fluidized region above. Provided such a transition is passive, its modelling as an interface is justified, and the theory bears this out. The static particle bed region is taken as a particle store, with the packed particles in static equilibrium (without strong cohesion). For the specific problem addressed in §6, namely the helicopter cloud problem, the continuum length scale $l$ will be taken as the geometric dimension of the helicopter blade length. The hydrodynamic effect of the helicopter is modelled as a fluid dipole, aligned with the axis of blade rotation and located at the helicopter blade rotation point, to represent the fluid down-draft, together with a line-vortex through the axis of rotation of the blades, to represent the fluid swirl.
@@ 3.2 The interfacial layer
which is the Magnus force induced on an individual rolling particle grain as it moves along the interfacial layer. We anticipate that in this incipiently fluidized interfacial layer this Magnus force dominates over the Saffman lift force due to the local shear in the fluid flow. The function $\Phi(E_{I})$ is anticipated to be, in general, monotone increasing in $E_{I}$, with $\Phi(E_{s})<1$ and $\Phi(1)=1$. Without detailed experimental evidence, we adopt the simple linear form approximation
@@
In (33) the local voidage in the interfacial layer $E_{I}$ is related to the local tangential fluid velocity in the interfacial layer, $\mathbf{u}_{I}$, whilst (32) relates the local tangential particle velocity in the interfacial layer $\mathbf{v}_{I}$ to the local tangential fluid velocity in the interfacial layer $\mathbf{u}_{I}$, and the form is similar to a mobile bed-load sediment transport function (see, for example, Bagnold (1941) for the classical discussion of bed-load sediment transport formula in a wind-blown sand environment). It should also be noted that throughout the above determination of the overall force balance on particles in the small continuum (containing many particles) element across the interfacial layer, we have taken the primary single particle representative forces for drag, friction, shear and lift to be as simple as possible, whilst retaining their principal features, and this is sufficient for our present purpose. However it is recognised that, if required, these basic forms could be refined, and thus improved, by introducing higher order contributions, using, for example, the recent results of Smith & Palmer (2019), Palmer & Smith (2020), Palmer & Smith (2021) and Jolley & Smith (2022), amongst others, arising from detailed studies of the motion of small free bodies located in the thin viscous boundary layer next to a wall, in high and moderate Reynolds number flows. We next consider the motion of the interfacial layer.
@@ 3.3 Interfacial layer motion
Figure 3: A qualitative schematic diagram representing a small element as a vertical slice through the static bed and the interfacial layer, with an illustration of the notation used in analysing this element.
@@ 5 The dimensionless problem
$\bar{R}_{f}E(\mathbf{u}_{t}+(\mathbf{u}\cdot\nabla)\mathbf{u})=-\nabla p-GE\mathbf{k}-(1-E)\bar{\beta}(E)(\mathbf{u}-\mathbf{v}),$ (53)
$\rho\bar{R}_{f}(1-E)(\mathbf{v}_{t}+(\mathbf{v}\cdot\nabla)\mathbf{v})=\alpha\nabla E-\rho G(1-E)\mathbf{k}+(1-E)\bar{\beta}(E)(\mathbf{u}-\mathbf{v}),$ (54)
@@
on $z=\xi$, with $\mathbf{r}_{h}\in\mathbb{R}^{2}$, $t>0$. Here we have introduced the dimensionless parameters $\rho=\rho_{s}/\rho_{f}$, giving the density ratio of the particulate material to the fluid material, and
$\bar{R}_{f}=\epsilon^{2}R_{f}$
where $\epsilon=a/l$ is the ratio of the microscopic scale particle radius and the macroscopic continuum length scale, and in general can be considered very small, whilst the fluid phase Reynold’s number, $R_{f}=\rho_{f}u_{s}l/\mu$, based on the macroscopic length scale l, is expected to be large, justifying our neglect of fluid phase macroscopic viscous terms in (53). Correspondingly $\bar{R}_{f}=\epsilon^{2}R_{f}$ measures the ratio of the inertia terms to the drag terms in the momentum equations (53) and (54). In what follows we will restrict attention to the situation when the macroscopic Reynolds number is large, but the ratio of the microscopic to macroscopic length scales is sufficiently small so that
$1\ll R_{f}\ll\epsilon^{-2},$ (61)
and hence $\epsilon^{2}\ll\bar{R}_{f}\ll 1$. As a consequence of this, in the fluid phase macroscopic momentum balance, inertia will play a secondary role. This is not uncommon in two-phase gas/solid particle flow, as seen in models of gas fluidized beds (see, for example, Needham & Merkin (1983)). In addition $G=\rho_{f}ga^{2}/\mu u_{s}$ measures the ratio of the gravity terms to the drag terms in the momentum equations (53) and (54), whilst $\alpha=p_{0}a^{2}/l\mu u_{s}$ measures the ratio of the particle-particle pressure term to the drag term in the momentum equation (54). Finally, $\gamma=3u_{s}^{2}\rho_{f}/ag\rho_{s}$ measures the ratio of the lift force to the gravity force acting on the particles in the interfacial layer. For convenience, we have also written
@@
$\rho\gg 1,\quad\bar{R}_{f}\ll\frac{\alpha}{\rho}\ll 1,\quad G\ll\frac{\alpha}{\rho}\ll 1.$
@@
$(2-E_{s}-E)\nabla\phi\cdot\hat{\mathbf{n}}+\alpha\nabla H(E)\cdot\hat{\mathbf{n}}=0,$ (76)
$\xi_{t}=\nabla\phi\cdot\hat{\mathbf{n}},$ (77)
@@
$\phi=\bar{\phi}+\mbox{{O}}(\alpha),$ (78)
with $\bar{\phi}=\mbox{{O}}(1)$ as $\alpha\rightarrow 0$, whilst, from (76) and (77) we may write
$\xi=\alpha\bar{\xi}+\mbox{{O}}(\alpha^{2}),$ (79)
with $\bar{\xi}=\mbox{{O}}(1)$ as $\alpha\rightarrow 0$, from which it follows that $\hat{\mathbf{n}}=\mathbf{k}+\mbox{{O}}(\alpha)$ as $\alpha\rightarrow 0$. On substituting from (78) and (79) into (70) and (76), the leading order problem for $\bar{\phi}$ decouples, and is given by
@@
With $E=\mbox{{O}}(1)$ as $\alpha\rightarrow 0$, the remaining problem for the voidage $E$ is given, from (71) and (75), as
@@ 6 The helicopter cloud problem
$\nabla\bar{\phi}=\frac{\Omega}{2\pi R}\hat{\boldsymbol{\theta}}+\mbox{{O}}(1),\quad\mbox{as }R\rightarrow 0\mbox{ uniformly in }\bar{D}_{c}.$ (85)
Here $(R,\theta,z)$ are the usual cylindrical polar coordinates in relation to the Cartesian coordinates $(x,y,z)$ which were introduced earlier, with unit vectors $\hat{\mathbf{R}}$, $\hat{\boldsymbol{\theta}}$ and $\mathbf{k}$. In (84), $\phi_{d}$ represents the harmonic half-space dipole, given by
@@
Throughout the above $\hat{\mathbf{N}}$ is the unit outward normal on the cylindrical shell. It remains to consider the far field boundary conditions at large distances from the cylindrical shell surface $\partial C$. We require
@@
$\nabla\bar{\phi}=\frac{\Omega}{2\pi R}\hat{\boldsymbol{\theta}}+\mbox{{O}}(1),\quad\mbox{as }R\rightarrow 0\mbox{ uniformly in }\bar{D}_{c}.$
@@
$\nabla\bar{\phi}(R,z)=a(R,z)\hat{\mathbf{R}}+\frac{\Omega}{2\pi R}\hat{\boldsymbol{\theta}}+b(R,z)\mathbf{k},$ (95)
@@
$E(\mathbf{r})=1+\mbox{{O}}(|\mathbf{r}|^{-2})\mbox{ as }|\mathbf{r}|\rightarrow\infty\mbox{ uniformly in }\bar{D}_{c}.$ (101)
@@
$H(E)=\frac{2}{9}(E-E_{s})+\mbox{{O}}(\delta\beta_{0}),$ (104)
as $\delta\beta_{0}\rightarrow 0$ uniformly for $E\in[E_{s},1]$. The complete form of [BVP] when written in terms of the cylindrical polar coordinates $(R,\theta,z)$ is presented in Appendix A, and its domain $\omega$ is illustrated in figure 6.
@@
Before proceeding further with [BVP], we must consider the detailed form of $g:[0,\infty)\to\mathbb{R}$, as defined in (100). First we examine $|\nabla_{h}\bar{\phi}(X,0)|^{2}$ for $X\in(0,\infty)$, the form of which depends upon the two parameters $\Omega$ and $z_{d}$. We begin by observing that
@@
We now consider the parameters in [BVP]. There are six dimensionless parameters, namely, $E_{s}$, $\alpha$, $\gamma$, $\Omega$, $z_{d}$ and $\Delta$. We can give order of magnitude estimates of these parameters for the helicopter cloud problem using typical helicopter parameters as recorded, for example, in the papers by Wachspress et al. (2008); Govindarajan et al. (2013); Porcù et al. (2020); Tan et al. (2021, 2022a, 2022b); Li et al. (2023); Lin et al. (2024) together with the standard properties of air and sand, to approximate $l,~{}M_{s},~{}\gamma_{s}$. The parameter $\Delta$ measures the ratio of the helicopter rotor core radius to the rotor blade radius, and typically $\Delta\approx 10^{-2}$. The parameter $z_{d}$ measures the ratio of the hovering height of the helicopter rotor to the rotor blade radius (recall figure 5), and typically $z_{d}\approx 10^{-1}$$10^{1}$. For the fluid the ratio of the hovering helicopter induced swirl velocity to the induced downwash velocity is measured by $\Omega$, and using typical values extracted from the above papers, we have $\Omega\approx 10^{-1}$$10^{0}$. The parameter $E_{s}$ is the particle packing voidage, and for sand in air this is typically of magnitude $10^{-2}$. The ratio of the lift force per unit volume to the gravity force per unit volume on particles in the interfacial layer is given by $\gamma$, and again using the typical values extracted from the above papers, we have $\gamma\approx 10^{2}$$10^{3}$. Finally, $\alpha$ measures the ratio of particle collisional pressure to drag in the particle flow phase and we tentatively estimate, for sand fluidized in air, that $\alpha\approx 10^{-1}$. With these estimates of all dimensionless parameters appearing in [BVP], in the next section we consider the numerical solution to [BVP].
7 Numerical examples
The accurate and efficient numerical solution of [BVP] requires a degree of nontrivial consideration, and as such, we present the technical details and assessment in Appendix B, where we also establish convergence of our numerical scheme (see example 5). Here, we now consider how the solution behaviour depends on our parameter choices. In examples 13, we fix
$E_{s}=0.01,\quad\alpha=0.1,\quad\Delta=0.04,$
and consider the nine possible combinations of
$\gamma=1000,500,100,\quad\Omega=1,0.5,0.1$
(with $\gamma=500$, $\Omega=0.5$, corresponding to example 5). As described in Appendix B, the accuracy of our numerical scheme depends on a number of discretisation parameters. From our assessment there of the scheme, we here choose the parameters $N=8$, $L_{R}=32N$, $L_{z}=16N$ (compared to $L_{R}=12N$ and $L_{z}=6N$ in example 5), and $\epsilon=1/(10N^{2})$, in order to ensure that the computational domain is large enough that all relevant solution behaviour will be captured across this wide range of values of $\gamma$ and $\Omega$. These parameter choices lead to the total number of degrees of freedom in our numerical solution being $\mbox{DOF}=2317056$ for $z_{d}=0.5$, $\mbox{DOF}=2481536$ for $z_{d}=0.9$ and $\mbox{DOF}=2646016$ for $z_{d}=1.3$, for each of examples 13. In each of the figures below, we truncate the computational domains where appropriate for presentational purposes, noting that the solution satisfies $E\approx 1$ everywhere outside the plotted range - note that all plots are over the same range within each figure, but that the plotted range varies between examples 13 ($R\in[0,6]$, $z\in[0,5]$ for example 1; $R\in[0,5]$, $z\in[0,4]$ for example 2; $R\in[0,3]$, $z\in[0,2]$ for example 3), though the computational domain for any given value of $z_{d}$ is identical for each example.
For all parameter choices in examples 1 and 2, and all in examples 3 except one (detailed below) we are either in case (a) or case (b)(i). This is illustrated for the case $\gamma=500$, $\Omega=0.5$, and $z_{d}=0.9$ or $z_{d}=0.5$ (as in example 5) in figure 8: for $z_{d}=0.9$, $\Omega=0.5>\Omega_{c}(0.9)$, and hence case (a) holds; for $z_{d}=0.5$, $\Omega=0.5<\Omega_{c}(0.5)$, and hence case (b) holds, however in this case both the local maximum and the local minimum take values greater than $(\gamma E_{s})^{-1}$ and $\gamma^{-1}$, hence case (b)(i) holds. The boundary data $g$ (shown in the lower part of figure 8) is qualitatively comparable in each of these cases.
Example 1 (Large $\gamma$)
Example 2 (Medium $\gamma$)
Example 3 (Small $\gamma$)
Again, we see more sand in the air (lower values of $E$) corresponding to larger values of $\Omega$, and to lower values of $z_{d}$. Comparing to figures 9 and 10, we see less sand in the air for the lower value of $\gamma$. Note that the plotted range is smaller in figure 11 than for both figure 10 and figure 9, though the computational domain is identical in each case. As for examples 1 and 2, for $\Omega=1$, and for $\Omega=0.5$ and $z_{d}=0.9,1.3$, the boundary data fits case (a), whilst for $\Omega=0.5$ and $z_{d}=0.5$, and for $\Omega=0.1$ and $z_{d}=0.9,1.3$ the boundary data fits case (b)(i), with similar qualitative solution behaviour. However, for $\Omega=0.1$ and $z_{d}=0.5$ we are now in case (b)(iii), as illustrated in figure 12. In this case, we see that the local maximum is greater than $(\gamma E_{s})^{-1}$, whilst the local minimum is less than $(\gamma E_{s})^{-1}$. As a result, the boundary data $g$ has a small hump near $X\approx 0.05$, though it is hard to see any discernible effect from this on the qualitative solution behaviour (bottom right plot of figure 11).
Example 4 (Two-plume example)
In our final example, we choose
$E_{s}=0.01,\quad\Delta=0.04,\quad\gamma=70,\quad\Omega=0.1,$
Figure 14: Voidage field $E$, example 4, $\alpha=0.1$, $z_{d}\in[0.4,1.1]$.
Figure 15: Voidage field $E$, example 4, $\alpha=0.02$, $z_{d}\in[0.4,1.1]$.
8 Conclusion
To illustrate the applicability of this theory we have formulated it in relation to a simple model of the helicopter cloud problem, leading to a fully determined nonlinear elliptic boundary value problem for the voidage field throughout the fluidized region. This provides a rational and efficient approach to computing the voidage field for this application. Careful numerical solution via a finite difference approximation scheme has demonstrated through a series of experiments how the qualitative behaviour of the voidage field varies as key parameters change, in a way that is entirely consistent with our understanding of the underlying physical processes. It may be of interest to look at this example in relation to controlled scale model experiments in order to specifically determine the qualitative and quantitative accuracy of the rational model that has been developed in this paper, but such considerations are left to future work. However, we do note that qualitative structural agreement between the high speed photography presented in Rovere (2022), and particularly relating to Figure 2.6 therein, and the theoretical figures presented here in §7, is very encouraging.